In the experiment diagrammed below, the grating ha
Answer #1

λ = 530 nm
d = 1/930 = 1.075 * 10^-5 m


7.3 * 10^-2 = λ*D/d [4 - 0]
7.3 * 10^-2 = (530 * 10^-9 * D * 4)/(1.075 * 10^-5)
D = 0.3703 m
D = 37.03 cm
Correct option :- (a)